Each equal side of an isosceles triangle is 3.5 inches longer than the base. The perimeter of the triangle is 2 feet. Find the length of the base.
Answer by ewatrrr(24785)

b + (b+3.5in) + (b+3.5in) = 24in
3b = 24-7
b = 17/3 0r 5 2/3 in